Find the value of the discriminant and determine the number of real-number solutions for this quadratic equation. y^2+3y-4=0
[0,one] [25,two] [30,one]
Can someone tell me how i find out which answer to pick and why

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Find the value of the discriminant and determine the number of real-number solutions for this quadratic equation. y^2+3y-4=0
[0,one] [25,two] [30,one]
-------------------
y^2+3y-4=0
It's in the form ax^2 + bx + c = 0
a = 1, b = 3, c = -4
---
The Disc is b^2 - 4ac
Disc = 3^2 - 4*1*(-4) = 9 +16 = 25
========================
Disc < 0 --> no real roots
Disc = 0 --> 2 real roots that are equal
Disc > 0 --> 2 real roots that are different

Answer by ikleyn(52778)

.
See the lessons
    - Introduction into Quadratic Equations
    - PROOF of quadratic formula by completing the square
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Quadratic equations".